let F(x,y) = <2x+yz,xz-2y,3z^2+xz> find the potential function.
Answer #1

28-2 38 +23 Given f(x, y) = (2x+43, ar-29, 38² +*8) we have to find potential ve fiuretion fox = 2x+y z fox = 3 3 +83 fy = 28